# PropMolFlow: Property-guided Molecule Generation with Geometry-Complete Flow Matching

### v3

@@ Results
Figure 2: Chemical validity and sampling efficiency of PropMolFlow against five baseline models. a, Molecule stability. b, RDKit validity. c, Uniqueness. d, PoseBusters validity. e, Closed-shell ratio. f, Sampling time for 10k molecules. Mean $\pm$ standard deviations are reported across six molecule properties. Structural validity metrics are shown in the value range between 0.4 and 1, while sampling times are in the unit of minutes.
Besides the ability to generate molecules with target properties, the structural validity and speed of generation are also critical in ensuring the success and robustness of chemical discovery with generative models. We further evaluated the performance of PropMolFlow by assessing the structural validity of the generated molecules and the sampling efficiency. Structural validity refers to whether the generated molecules conform to basic chemical rules. This was measured using several metrics, including atomic stability, molecule stability, RDKit validity, ‘uniqueness and validity’, PoseBusters validity, and closed-shell ratio. Atom stability is given by the proportion of atoms with correct valency. To account for atomic charges, an atom is considered to have the correct valency if its formal charge balances its explicit valency, as defined in RDKit [33]. For instance, a nitrogen atom should carry a +1 formal charge if it has a valency of 4. Molecule stability defines the proportion of molecules in which all atoms are stable and the molecule is charge neutral (zero net formal charges). RDKit validity refers to the ratios of molecules that pass RDKit’s sanitization check. ‘Uniqueness and validity’ is defined as the proportion of molecules that are RDKit valid and unique in their SMILES representation. PoseBusters validity corresponds to the fraction of molecules that pass many de novo chemical and structural validation tests, such as all atom connectivity, valid bond lengths and angles, and absence of internal clashes [34]. A molecule is closed-shell valid if it has an even number of valence electrons. Since all molecules in QM9 are closed-shell, any generated open-shell species are considered invalid. Sampling efficiency is given by the wall-clock time required to generate 10,000 molecules. Computational settings for sampling are provided in Methods.
Across nearly all stability and validity benchmarks, PropMolFlow consistently outperforms baseline modes (Figure 2a–e). The only exception is the closed-shell ratio, where it slightly underperforms JODO—a diffusion model that omits atomic formal charges and adds additional bond features such as aromaticity existence. Moreover, PropMolFlow achieves high chemical validity with much faster sampling efficiency. Flow matching is normally much faster than diffusion models due to its shorter, deterministic probability path (Eq. (6) and (8)) and optimal transport (Supplementary Section LABEL:sec:ot[35, 21]. PropMolFlow requires only 100 time steps compared to the 1000 diffusion time steps, and achieves a speedup for sampling of at least 8$\times$ over diffusion-based models and it is nearly 2$\times$ faster than EquiFM (Figure 2f).
When evaluating metrics on a per-property basis, PropMolFlow consistently achieves the highest molecular stability and validity across all properties (Supplementary Tables LABEL:tab:propmolflow_metric_by_property-LABEL:tab:jodo_metric_by_property). Across properties, PropMolFlow’s metrics follow the order: $\mu\approx\Delta\epsilon>\epsilon_{\mathrm{HOMO}}>\epsilon_{\mathrm{LUMO}}>\alpha>C_{v}$. This trend generally holds for the baseline models as well, with the exception of EEGSDE, suggesting a more complex structure–property relationship for $\alpha$ and $C_{v}$. This complexity may be due to their strong dependence on molecular geometry and the limitation of the training data, which includes only fully relaxed molecules. Across metrics, molecular stability and “Validity & Uniqueness” exhibit over 2% variations across properties, whereas molecular validity and PoseBusters validity vary by less than 1%. Besides, all methods in this study generate a non-negligible fraction of open-shell molecules: on average 1.6% for JODO and at least 4% for the other methods. This highlights a previously overlooked issue in generative molecular modeling: the inability to consistently respect the valence electron constraints inherent to the training data. Addressing this limitation is essential for improving the chemical validity of generated molecules especially when atomic formal charges are considered in molecular graphs, and future work should improve upon this direction—perhaps by explicitly including closed-shell constraints, as we do here by adding bond orders into molecular graphs—to ensure better alignment with the reference distribution.
Property predictor as DFT proxy reveals systematic inductive bias
Figure 3: Performance of GVP property predictors for PropMolFlow-generated molecules without and with relaxation. a, Comparison between Target, DFT and GVP shows the reliability of GVP to evaluate the MAE metrics commonly used for property-guided generation. Comparison between GVP and GVP-R, and between DFT and DFT-R shows the structural dependence of GVP-predicted and DFT-predicted properties, respectively. Comparison between GVP and DFT with and without relaxation shows the reliability of GVP in capturing the ground-truth DFT values over raw and relaxed structures. Target is the input target property values used by PropMolFlow models to generate models. ‘DFT’ and ‘GVP’ are respective DFT-computed and GVP-evaluated property values on directly generated raw molecules. ‘DFT-R’ and ‘GVP-R’ refer to values evaluated on DFT-relaxed molecules. b, Pairwise comparison between Target, DFT and GVP on raw molecules. c, GVP versus DFT for both raw and DFT-relaxed molecules. d, Structure variation and property dependence on the DFT relaxation. Molecules with the highest RMSD for each property and their DFT and GVP values are listed as examples. In molecule representation, gray, red, blue, and white colors indicate C, O, N, and H, respectively. Property values for $\alpha$, $\Delta\epsilon$, and $C_{v}$ are in units of Bohr3, eV, and cal/(mol $\cdot$ K), respectively.
A separate surrogate model is needed to evaluate the properties of generated molecules. Typically, this surrogate model shares the same architecture as the generative neural network model [7]. However, using models from the same family may introduce inductive biases, making it necessary to validate results using external property predictors, such as DFT. To this end, we performed DFT calculations on 1000 molecules selected from the 10,000 structures generated for each property in the ID task. For directly generated molecules, multi-step filtering is essential to ensure that the structures are chemically valid and suitable for downstream applications. We filtered the 1000 molecules by five criteria: Molecule stability, RDKit validity, PoseBusters validity, closed-shell validity, and multi-fragment check. The multi-fragment check filters out molecules that with more than one disconnected fragment. For molecules passing all filters, property values are evaluated using both a GVP property predictor and DFT calculations. We also filtered out molecules whose DFT calculations did not converge, resulting in 90% of the molecules remaining for the subsequent analyses (Table LABEL:tab:no_sp_dft). To evaluate the reliability of GVP property predictor and the impact of structural relaxation, we conducted a comprehensive comparison among target property values (Target), DFT-calculated properties (DFT), and GVP predictions (GVP)—using both directly generated and DFT-relaxed molecules (Figure 3a). We chose three example properties, $\alpha$, $\Delta\epsilon$ and $C_{v}$, for the following analyses because the results for the other three properties are similar to those of $\Delta\epsilon$ (Supplemental Figure LABEL:fig:si_target_dft_gvp). Example orders of pairwise MAE distances for $\alpha$ are also provided in Figure 3a.
We first assess the reliability of GVP on the initially generated molecules (Figure 3b). The MAEs between the target values and DFT (Target vs. DFT) are comparable to those betwen the target and GVP predictions (Target vs. GVP) for $\alpha$ and $\Delta\epsilon$, suggesting that GVP property predictors are statistically reliable for estimating properties of directly generated molecules. However, GVP consistently underestimates the MAEs, revealing an inductive bias in its predictions. Notably, although the MAEs of GVP and DFT relative to the target are similar, the MAEs between GVP and DFT themselves are significantly larger—for example, 0.91 vs. 0.1 Bohr3 for $\alpha$, and 204 vs. 43 meV for $\Delta\epsilon$. For $C_{v}$, the MAE between the target values and DFT is much larger than that between the target values and GVP. However, this discrepancy is not due to poor reliability of the GVP property predictors, but rather stems from an issue with DFT-calculated heat capacities ($C_{v}$) on unrelaxed structures. DFT-computed $C_{v}$ values depend on vibrational frequencies, which are highly sensitive to molecular geometry. To validate this analysis, we relaxed the structures, resulting in the DFT vs. Target MAE dropping from 1.68 to 0.68 cal/(mol $\cdot$ K), closely aligning with the MAE of 0.61 cal/(mol $\cdot$ K), predicted by GVP.
Evaluation on directly generated molecules without geometry optimization reflects the statistical quality of the learned structure–property relationships captured by both PropMolFlow models and the GVP predictors. However, from a chemical discovery perspective, it is essential to relax the structures and evaluate properties based on the relaxed chemical stable conformations that are more likely to occur in real-world experiments. To assess the reliability of GVP on relaxed structures, we compare its predictions with DFT-computed property values for both unrelaxed and relaxed molecules (Figure 3c). For $C_{v}$, a much closer agreement between DFT and GVP was observed on relaxed structures. In contrast, for $\alpha$, a notable offset between GVP and DFT appears when using relaxed structures, indicating that GVP is more reliable on unrelaxed structures. This can be attributed to the inductive bias of GVP property predictors—specifically, GVP tends to perform better on directly generated structures by PropMolFlow, as both models share the same architecture. This bias is further supported by the increased MAEs between GVP and DFT after relaxation for all properties except $C_{v}$ (Supplemental Figure LABEL:fig:id_mae_heatmap). For $\Delta\epsilon$, despite a slight broadening of the peak at lower $\Delta\epsilon$ values is observed, both DFT and GVP predictions remain largely unchanged, suggesting that this property has weak dependence on structural relaxation.
To further quantify the effect of relaxation, we calculated the symmetry-aware root mean square distances (RMSDs) between atomic positions of raw structures and DFT-relaxed structures. Given the same chemical compositions and element-type ordering, the RMSD between any two molecules (M0 and M1) is the minimum distance between these two structures considering the translational and rotational symmetry operations, which can be expressed as:
Where $SE(3)$ represents the group of operations that respect the translational and rotational symmetry. The RMSD distributions for $\alpha$, $\Delta\epsilon$ and $C_{v}$ are showed in Figure 3d. All RMSD distributions are highly left-skewed and concentrated around 0 Å, meaning the majority of PropMolFlow-generated structures are close to their relaxed, stable counterparts. To estimate the property sensitivity to structural relaxation, we selected the structures whose RMSDs are larger than 0.03 Å—RMSDs lower than this threshold are considered very close to the relaxed structures, and in Eq. (2), we define the normalized structure sensitivity of DFT-calculated properties to allow comparisons between different properties.
where $\delta q_{\rm{DFT}}$ is the change of DFT-calculated property upon relaxation, and the normalizer ($q_{0.99}-q_{0.01}$) is the difference between the 99th quantile and the 1th quantile of the property distribution of QM9 training data. This range is chosen over the min-max range because it is robust to outliers. Despite similar RMSD distributions across the three properties, the differences are more pronounced in the structural dependence of DFT-calculated properties, following the order of $C_{v}>\alpha>\Delta\epsilon$. On the right panel of Figure 3d, molecule configurations with the highest RMSDs are shown, and their RMSD, taget input, DFT-calculated, GVP-predicted property values are listed to rationalize the relationship between GVP, DFT and Target values with and without relaxation. DFT property values for relaxed structures can be either closer or farther to the target input values. For these molecules with large structural changes, the GVP predictions are nearly unaltered, showing a much weaker structural dependence compared to that of DFT. Although DFT property changes are more significant than GVP upon relaxation for the three extreme cases, GVP predictions are close to DFT-calculated values for both raw and relaxed molecules, offering reliable evaluation of PropMolFlow models and robust guidance towards chemical discovery.
Interpolated structures follow chemically intuitive trends
To assess whether PropMolFlow models have learned a continuous and smooth structure–property relationship, we performed interpolation by varying the target property values while fixing the number of atoms in the generated molecules. For each property value, ten molecules were sampled, and the one that was chemically valid and whose DFT-calculated property most closely matched the target property value was selected. For each property, we show the minimum and maximum target values along with their DFT-calculated values below the molecular configurations. The target property value, DFT-calculated value, and GVP predictions for each configuration are provided in Supplemental Table LABEL:tab:interpolation.
Figure 4: PropMolFlow generated molecules when interpolating among different property values while fixing the number of atoms at 19 for sampling. The minimum and maximum target properties in red and corresponding DFT-calculated properties in blue are provided below the configurations. All molecules chosen here pass the five filtering criteria and have DFT values closest to the target properties. C, H, O, and N are in gray, white, red, and blue colors, respectively. Property units are provided in the square brackets under each property symbol.
@@
The interpolated structures generated by PropMolFlow exhibit smooth and chemically intuitive transitions as property values vary, demonstrating that the model has learned meaningful structure—property relationships. These results highlight the model’s ability to generate property-guided molecules that align with established chemical principles.
@@
Figure 5: Performance of PropMolFlow in OOD generation. a, Distribution of DFT calculated and GVP predicted values for PropMolFlow generated molecules, and the property distribution of QM9 is also included. The vertical black dashed line in histograms represents the target property value $q_{0.99}$. Curves on top of histograms are fitted with a kernel density estimation. b, Three example molecules that do not exist in QM9 but are found in a larger PubChem dataset are included in the left panel. Numbers below the configurations are DFT calculated property values on raw molecules generated by PropMolFlow models. C, H, O, N, and F are in gray, white, red, blue, and yellow colors, respectively. Property values for $\alpha$, $\Delta\epsilon$, and $C_{v}$ are in units of Bohr3, eV, and cal/(mol $\cdot$ K), respectively.
PropMolFlow pushes molecule generation toward out-of-distribution property space
Although ID task is typically used to evaluate property-guided generative models, it is also of practical interest to test molecular generation for properties that are underrepresented in the training dataset. To this end, we propose an ‘out-of-distribution’ (OOD) task designed to assess the model’s ability to generate molecules conditioned on property values that are rare or sparsely represented in the original training distribution. In this task, we chose the 99th quantile of the property distribution of QM9 training data as the target property value. The target value for each property is provided in Supplemental Table LABEL:tab:prop_choice. We sampled 1,000 structures conditioned on the chosen property value using the models with top-performing embedding methods (Supplemental Table LABEL:tab:best_model). These generated molecules then pass through the same multi-step filtering procedure used in the ID task, and their properties are evaluated by both GVP and DFT.
Here we compare the property distribution of generated molecules against that of the QM9 training data for $\alpha$, $\Delta\epsilon$ and $C_{v}$ (Figure 5a). Results for other properties are provided in Supplemental Figure LABEL:fig:si_ood. The DFT results demonstrate that the OOD generation successfully shifts the property distribution of generated molecules towards the specified target values (indicated by the dashed line), and the DFT-predicted distribution is nearly perfectly centered on the target property value for the polarizability ($\alpha$). However, for $\Delta\epsilon$ and $C_{v}$, the DFT-predicted distribution is off the target by a small margin. For example, both the GVP and DFT distributions for the gap are centered around 8.95 eV, which is 0.41 eV away from the target value of 9.36 eV. This deviation is likely due to the scarcity of training data in the corresponding region of property space. One should note the DFT issue for $C_{v}$ on unrelaxed structures which can be largely addressed by performing relaxation. While property guidance effectively steers the generative model toward the desired values, the model’s capacity to extrapolate is ultimately constrained by the coverage of the training distribution. Other guidance approaches, such as classifier-free guidance [36], can adjust the weights for the property conditioning, potentially achieving better alignment with target property values, as demonstrated in the conditional generation of inorganic crystal materials [37]. Moreover, the GVP distributions are almost overlapped with the DFT ones, showing desired extrapolation capability of the GVP property predictors. To further evaluate the novelty of generated molecules, Figure 5b presents three example molecules that are not part of the QM9 dataset but are found in the larger PubChem database [38]. These molecules exhibit the DFT properties close to their target input values. This demonstrates PropMolFlow’s ability to generate chemically valid molecules with desired target properties that lie outside the original training domain.
Conclusion
In this work, we introduce PropMolFlow, a property-guided molecule generation framework based on the SOTA SE(3) flow matching. Our approach outperforms existing diffusion and flow matching models in the conditional generation setting, offering improved property alignment, structural validity and sampling efficiency across many molecular properties. We conduct a systematic evaluation of property embedding strategies and identify optimal configurations that depend on both the target property and the task regime—whether ID or OOD. To enable a more comprehensive evaluation, we propose two new metrics and one new task: The closed-shell ratio and revised atomic stability and molecular stability to account for charged species, which quantify how well the generated molecules align with the training data in terms of valency–charge configurations, and an OOD task, designed to evaluate controllability in sparsely represented regions of property space.
To validate the fidelity of generated molecules and reliability of property predictors, we perform extensive DFT calculations. PropMolFlow generates molecules that are close to the stable, relaxed counterparts. Generally, GVP property predictors provide reliable approximations of the true DFT MAEs against target values despite the relative large difference between GVP and DFT. We demonstrated the inductive bias of property predictors, which lead to a systematic underestimation of true MAEs and its partial failure when predicting on unseen relaxed structures. Additionally, we curated a DFT dataset of generated molecules, which is made available through the ColabFit Exchange [31, 39]. While this work focuses on small molecules from the QM9 dataset, PropMolFlow is readily extendable to larger molecular datasets with more atoms, as well as to materials datasets, as long as property labels are available.

### v4

@@ Introduction
Structural validity and inference speeds are critical in success of chemical discovery with generative models. Structural validity refers to whether generated molecules conform to chemical rules. This was measured using five metrics, including atomic stability, molecule stability, RDKit validity [rdkit], ‘uniqueness and validity’, PoseBusters validity [buttenschoen2024], and closed-shell ratio. Inference speeds are evaluated by the wall-clock time for generating 10,000 molecules. Details of structural metrics and computational settings for sampling are provided in respective ‘Evaluation metrics’ and ‘Computational settings’ section.
@@
[…]
Systematic inductive bias of property predictors
Figure 3: Performance of GVP property predictors without and with DFT relaxation. a, Comparison between Target, DFT and GVP shows the reliability of GVP in evaluating MAE metrics commonly used for property-guided generation. Comparison between GVP and GVP-R, and between DFT and DFT-R shows the structural dependence of GVP-predicted and DFT-predicted properties, respectively. Comparison between GVP and DFT with and without relaxation shows the reliability of GVP in capturing ground-truth DFT values for both raw and relaxed structures. ‘Target’, ‘DFT’ and ‘GVP’ denote input, DFT-calculated and GVP-predicted property values on raw molecules, respectively. ‘DFT-R’ and ‘GVP-R’ refer to values evaluated on DFT-relaxed molecules. The $d$ indicates the MAE distances between two property-value vectors. b, Pairwise comparison between Target, DFT and GVP on raw molecules. c, GVP versus DFT for both raw and DFT-relaxed molecules. d, Root mean squared distances (RMSDs) and normalized property sensitivity due to DFT relaxation. Molecules with the highest RMSDs for each property and their corresponding DFT and GVP values are shown. In molecular representations, gray, red, blue and white indicate C, O, N and H atoms, respectively. Property values for $\alpha$, $\Delta\epsilon$ and $C_{v}$ are in units of Bohr3, eV and cal mol-1 K-1, respectively.
A property predictor is needed to evaluate the generated molecules. Typically, this predictor shares the same architecture as the generative model, which may introduce inductive biases in predictions. In this work, DFT calculations were carried out on selected and filtered molecules and compared to input (‘Target’) and GVP-predicted values, without and with (‘-R’) structural relaxations. For example, we show orders of pairwise MAE distances for $\alpha$ in Fig. 3a. We show results for $\alpha$, $\Delta\epsilon$ and $C_{v}$ in Fig. 3b–d, and for the other properties in Supplementary Figure LABEL:fig:si_target_dft_gvp, which are similar to $\Delta\epsilon$. DFT settings are included in the ‘DFT’ section, the filtering procedure are provided in the ‘Molecule filtering’ section, and numbers of filtered molecules are listed in Supplementary Table LABEL:tab:no_sp_dft.
First, we examine GVP on the raw generated molecules (Fig. 3b). MAEs of “Target vs. DFT” are comparable to “Target vs. GVP” for $\alpha$ and $\Delta\epsilon$, suggesting that GVP predictors are reliable for estimating “Target vs. DFT” MAEs. However, GVP consistently underestimates the MAEs, revealing an inductive bias. For $C_{v}$, the “Target vs. DFT” MAE is much larger than “Target vs. GVP”. This is because DFT-computed $C_{v}$ depends on vibrational frequencies, which are highly sensitive to geometry. Relaxation resolves this issue, reducing the MAE to 0.68 cal/(mol $\cdot$ K), still higher than, but closer to “Target vs. GVP” (0.61). For chemical discovery, it is also important to relax the structures. Fig. 3c shows a much closer agreement between DFT and GVP after relaxation for $C_{v}$, while a notable offset is observed for $\alpha$. It indicates that GVP is more reliable on unrelaxed structures, perhaps due to its inductive bias. This is also true for other properties (Supplementary Figure LABEL:fig:id_mae_heatmap).
To further quantify the relaxation effect, we calculated the root mean square distances (RMSDs) between atomic positions before and after relaxation. Fig. 3d shows that RMSD distributions are left-skewed toward 0 Å, meaning most unrelaxed structures are close to their relaxed counterparts. To estimate the property sensitivity to relaxation, we calculated the normalized property sensitivity to allow comparisons between different properties. Definitions of RMSDs and normalized property sensitivity are provided in the ‘Evaluation metrics’ section. Despite similar RMSD distributions, the differences are more pronounced in the property sensitivity, following the order of $C_{v}>\alpha>\Delta\epsilon$. Fig. 3d also shows configurations with the highest RMSDs, their target, DFT-calculated, and GVP-predicted values. DFT values with relaxation can be either closer or farther to Target. Despite high RMSDs, GVP values are nearly unaltered with relaxation, showing a much weaker structural dependence of GVP versus DFT. Overall, GVP predictions are close to DFT values for all cases, hence it can offer a statistically reliable evaluation of PropMolFlow models despite the inductive biases.
Interpolated structures aligned with chemical intuition
To assess whether PropMolFlow has learned a smooth structure–property relationship, we performed interpolation by varying target values while fixing atom counts of molecules at 19. For each value, ten molecules were sampled, and the one whose DFT-calculated property is closest to the input was selected. We show the configurations for selected molecules, and minimum and maximum target values of each property along with corresponding DFT values in Fig. 4. Numeric results for all configurations are provided in Supplementary Table LABEL:tab:interpolation.
Figure 4: Interpolation study by varying property values. The minimum and maximum target properties (red) and the corresponding DFT-calculated properties (blue) are shown below each configuration. All molecules shown pass the filtering criteria and have DFT values closest to the target properties. In molecular representations, gray, white, red and blue indicate C, H, O and N atoms, respectively. Property units are provided in square brackets under each property symbol.
[…]
Toward out-of-distribution generation
We propose an OOD task to test generation for underrepresented property values. The 99th quantile of the training data distribution was chosen as the target value and the atom counts were chosen according to the ‘Inference’ section. Specific target values are provided in Supplementary Table LABEL:tab:si_prop_val_stats. 1,000 molecules were sampled with top-performing models in this task (Supplementary Table LABEL:tab:best_model), and we filtered structures using the same procedure as that used in the ID task. We compare DFT and GVP property distributions against that of the QM9 training data (Fig. 5a). DFT distributions shift toward the targets (dashed vertical lines). The distribution mode is very close to the target for $\alpha$. For $\Delta\epsilon$ and $C_{v}$, DFT distribution modes are off the targets by a small margin, for example 0.41 eV for $\Delta\epsilon$ (8.95 versus target 9.36 eV). This deviation is likely due to the scarcity of training data in the corresponding property region; when conditioning on the 50th quantile (medians), modes are closer to the targets (Supplementary Figure LABEL:fig:si_q05_vs_q099). Moreover, GVP distributions are almost overlapped with DFT ones, suggesting their desired extrapolation performance.
To examine the novelty of generated molecules, Fig. 5b presents three generated molecules whose SMILES representations are not part of QM9 but are found in the larger PubChem database [kim2021]. We also evaluated the novelty of generated molecules by Tanimoto similarity to the training data. Fig. 5c indicates that 20–35% of the molecules exist in the training data. Using 0.8 as a cutoff, novelty ratios are 58–75% across all properties, demonstrating substantial exploration beyond the training data distribution. Results for other properties are provided in Supplementary Figure LABEL:fig:si_ood, and additional extrapolation to more extreme property values or atom counts can be found in respective Supplementary Figure LABEL:fig:si_various_values and Supplementary Figure LABEL:fig:si_ood_no_atoms.
@@
Figure 5: Toward out-of-distribution generation. a, Distribution of DFT-calculated and GVP-predicted property values for molecules generated by PropMolFlow; the property distribution of the QM9 training data is also shown. The vertical black dashed line in the histograms denotes the target property value $q_{0.99}$, corresponding to the 99th percentile of the training-data distribution. Curves overlaid on the histograms are kernel density estimation fits. b, Three example molecules absent from QM9 but present in a larger PubChem dataset are shown in the left panel. Numbers below the configurations indicate DFT-calculated property values for raw molecules generated by PropMolFlow. In molecular representations, gray, white, red, blue and yellow indicate C, H, O, N and F atoms, respectively. Property values for $\alpha$, $\Delta\epsilon$ and $C_{v}$ are in units of Bohr3, eV and cal mol-1 K-1, respectively. c, Maximum Tanimoto similarity between generated, filtered molecules and the training data computed using Morgan fingerprints. Dashed lines indicate the 0.8 similarity cutoff used to define novel molecules.
Discussion
While PropMolFlow can generate chemically valid and novel molecules with strong property alignment and fast inference, its ability to extrapolate could be further enhanced, potentially by integrating the model into an active-learning or reinforcement-learning framework. The current implementation is limited to the small-molecule QM9 dataset; extending PropMolFlow to larger molecular datasets and conditioning on properties of greater practical relevance represents an important future direction. In some cases, the property distribution of generated molecules exhibits deviations from target values. Property alignment could be improved by incorporating advanced guidance mechanisms with tunable guidance weights [ho2022, karras2024]. Moreover, despite faster inference than diffusion models, flow matching still requires multi-step integration . Additional efficiency gains may be possible by learning flow maps that support direct jumps along probability paths [boffi2025]. However, these approaches have shown effectiveness only on continuous data; learning flow maps for multi-modal 3D molecules remains a challenge.
Although extending PropMolFlow to multiple properties is conceptually straightforward, the optimal strategy for synthesizing different properties and for managing potential correlations remains unclear. Another challenge concerns conformational stability: PropMolFlow does not guarantee the generation of the most stable conformers. Injecting energies into de novo generation is an open question, though recent advances in conformer generation with fixed compositions, such as adjoint Schrödinger bridge sampling [liu2025], may provide useful insights.
Furthermore, several techniques, such as fake atoms and self-conditioning, were recently introduced in the unconditional FlowMol model [dunn2025-3] while this work was under development, and may further benefit the PropMolFlow conditional model. Despite these limitations and challenges, PropMolFlow demonstrates reliable generative modeling in chemical discovery and provides a strong foundation for future advancements.